OpenStudy (scotty):

what is 3/5 of 30? and how do you work this out?

OpenStudy (shadowfiend):

Whenever you say `some fraction OF some number', you can multiply the two together to get your number. So, in this case: 3/5 OF 30 means: \[\frac{3}{5}\times 30\]

OpenStudy (shadowfiend):

Can you do that multiplication?

OpenStudy (scotty):

No

OpenStudy (scotty):

do I do 3*30 and then 5*30?

OpenStudy (scotty):

i've found the answer!!!

OpenStudy (shadowfiend):

AWESOME!

OpenStudy (shadowfiend):

Hehe, sorry for abandoning you there for a minute. That was the equivalent of: \[\frac{3}{5}\times \frac{30}{1}\] So you can multiply the 3 across and keep the 5, then do the division.

OpenStudy (scotty):

is the answer 18?

OpenStudy (apples):

Yes, it is. \[\frac{3}{5} \cdot \frac{30}{1} = \frac{3 \cdot 30}{5 \cdot 1} = \frac{3 \cdot 6}{1} = \frac{18}{1} = 18\]